### v6

# Phase transition near the Big Bang in the lattice theory of gravity and some cosmological considerations

Abstract
Lattice regularization of the theory of gravity provides a new possibility for study of the Big Bang problem. Moreover, the very possibility of studying the state of the Universe at the highest temperatures appears. In the 4D lattice theories of gravity, the existence of a high-temperature phase is proved, which is characterized by the folding of space into a point and zeroing of the energy-momentum tensor. This fact leads to the hypothesis that there are at least two phases in the 4D lattice theory, and the main properties of the high-temperature phase are similar to those described above. The solution of the Einstein equation in the low temperature phase and the long-wave limit within the framework of the Friedmann paradigm with a finite bare cosmological constant is given. This solution is mathematically correct, since all local physical quantities (energy density including vacuum energy, etc.) on the lattice are finite. As a result, a solution is obtained that demonstrates an exponential growth of the cosmological scale factor $a(t)$ in the initial period of evolution (inflation phase of the Universe) and then passes into a power law ($a(t)\propto\sqrt{t}$). This solution also demonstrates that the energy-momentum tensor tends to zero when approaching the Big Bang point.
pacs:
04.60.Bc
I Introduction
In the works of the author Vergeles (2006, 2015, 2017a, 2017b, 2021a) a model of discrete gravity on a simplicial complex was studied. This model included pure gravity degrees of freedom (tetrads, connection), as well as a fermion Dirac field minimally coupled to gravity. In particular, in the work Vergeles (2021a) a discrete symmetry $\mathbb{Z}_{2}$ of the action called PT symmetry was described. This PT transformation changes the sign of all tetrads and mutually interchanges the Dirac variables with their Hermitian conjugates, that is, mutually interchanges particles and antiparticles.
In this work, we take the next step: we show that at ultrahigh temperatures the PT symmetry is not broken, but it is broken at sufficiently low temperatures. In the high-temperature phase, the mean values of the energy-momentum tensor of matter, as well as the mean values of the tetrads, are equal to zero. On the contrary, in the low-temperature phase these values are different from zero. Tetrad is an order parameter.
As is known, in the Minimum Standard Model (MSM) (considered within the framework of gravity) the number of fermion degrees of freedom is a multiple of the number of bosonic degrees of freedom. Therefore, we assume that some properties of the MSM near the point of the Big Bang can be described using the discrete gravity model being studied. Conversely, the following question is of interest to us: does the continuation of solutions to Einstein’s continuum equations in the vicinity of the Big Bang point contain some information about the properties of the high-temperature phase in lattice theory? To answer this question, we solve Einstein’s equations in the spirit of Friedmann’s model, but with a bare cosmological constant of positive sign. The positive cosmological constant compensates for the enormous energy density of negative-sign quantum zero-point oscillations of the vacuum. Here we make the following remarks: 1) The energy density of zero-point vacuum oscillations is negative, since fermion degrees of freedom predominate in the MSM. 2) We consider the continuum model of gravity with matter fields included as the long-wave limit of the lattice theory. Therefore, all quantities (energy density $\varepsilon$, bare cosmological constant $\Lambda_{0}$, etc.) are finite. In this case, Einstein’s equations and their solutions are mathematically correct. The solution found in this work shows a decrease in the energy density modulus $|\varepsilon|$ to zero when approaching the Big Bang point. Of course, in the vicinity of the Big Bang point, neither the equations nor the solution are correct due to the increase in the value of quantum fluctuations (see Section IV). However, the tendency towards a decrease in $|\varepsilon|$ within the framework of the reliability of the solution indicates the consistency of the lattice model and its long-wave limit. Let us emphasize that the decrease in $|\varepsilon|$ occurs in parallel with an increase in the density of matter.
The article is organized as follows.
Since different variants of simplicial theories of gravity are discussed, for ease of reading we define the variant studied here in Section II.
Section III proves the main result of the work: in the lattice theory of gravity, discrete PT symmetry is not broken at ultrahigh temperatures. But at low temperatures, spontaneous breaking of this symmetry occurs.
Section IV considers the solution of the Einstein equation within the Friedmann paradigm in the long-wave limit. The validity of the classical approach (without taking into account quantum fluctuations) at some distance from the Big Bang point is substantiated.
Section V provides some thermodynamic considerations.
This work is of a model nature.
II introduction to lattice theory of gravity
II.1 Lattice theory definition
To go further, it is necessary to formulate the variant of the theory of gravity on a lattice, which we will use. For a more detailed description of this theory, see Vergeles (2006, 2015, 2017a, 2017b, 2021a)
In this and next Sections, consideration is carried out for the Euclidean signature.
Denote by $\gamma^{a}$ $4\times 4$ Hermitian Dirac matrices, so that
$\gamma^{a}\gamma^{b}+\gamma^{b}\gamma^{a}=2\delta^{ab},\quad\sigma^{ab}\equiv\frac{1}{4}[\gamma^{a},\gamma^{b}],\quad a=1,2,3,4,\quad\gamma^{5}\equiv\gamma^{1}\gamma^{2}\gamma^{3}\gamma^{4}=(\gamma^{5})^{{\dagger}}.$ (1)
Consider the orientable abstract 4-dimensional simplicial complex $\mathop{{\mathfrak{K}}}\nolimits$. Suppose that any of its 4-simplexes belongs to such a finite (or infinite) sub-complex ${\mathop{{\mathfrak{K}}}\nolimits}^{\prime}\in\mathop{{\mathfrak{K}}}\nolimits$ which has a geometric realization in $\mathbb{R}^{4}$ topologically equivalent to a disk without cavities. The vertices are designated as $a_{{\cal V}}$, the indices ${{\cal V}}=1,2,\dots,\,{\mathop{{\mathfrak{N}}}\nolimits}^{(0)}\rightarrow\infty$ and ${{\cal W}}$ enumerate the vertices and 4-simplices, correspondingly. It is necessary to use the local enumeration of the vertices $a_{{\cal V}}$ attached to a given 4-simplex: the all five vertices of a 4-simplex with index ${{\cal W}}$ are enumerated as $a_{{{\cal V}}_{({{\cal W}})i}}$, $i=1,2,3,4,5$. The later notations with extra low index $({{\cal W}})$ indicate that the corresponding quantities belong to the 4-simplex with index ${{\cal W}}$. Of course, these same quantities also belong to another 4-simplex with index ${{\cal W}}^{\prime}$, and 4-simplexes with indices ${{\cal W}}$ and ${{\cal W}}^{\prime}$ must be adjacent. The Levi-Civita symbol with in pairs different indexes $\varepsilon_{{{\cal V}}_{({{\cal W}})1}{{\cal V}}_{({{\cal W}})2}{{\cal V}}_{({{\cal W}})3}{{\cal V}}_{({{\cal W}})4}{{\cal V}}_{({{\cal W}})5}}=\pm 1$ depending on whether the order of vertices $s^{4}_{{\cal W}}=a_{{{\cal V}}_{({{\cal W}})1}}a_{{{\cal V}}_{({{\cal W}})2}}a_{{{\cal V}}_{({{\cal W}})3}}a_{{{\cal V}}_{({{\cal W}})4}}a_{{{\cal V}}_{({{\cal W}})5}}$ defines the positive or negative orientation of 4-simplex $s^{4}_{{\cal W}}$. An element of the compact group $\mathop{\rm Spin}\nolimits(4)$ and an element of the Clifford algebra
$\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}=\Omega^{-1}_{{{\cal V}}_{2}{{\cal V}}_{1}}=\mathop{\rm exp}\nolimits\left(\omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\right)=\mathop{\rm exp}\nolimits\left(\frac{1}{2}\sigma^{ab}\omega^{ab}_{{{\cal V}}_{1}{{\cal V}}_{2}}\right)\in\mathop{\rm Spin}\nolimits(4),\quad\sigma^{ab}\equiv\frac{1}{4}[\gamma^{a}\gamma^{b}]$
$\hat{e}_{{{\cal V}}_{1}{{\cal V}}_{2}}\equiv e^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\gamma^{a}\equiv-\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\hat{e}_{{{\cal V}}_{2}{{\cal V}}_{1}}\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}^{-1},$
$|e_{{{\cal V}}_{1}{{\cal V}}_{2}}|<1,\quad|e_{{{\cal V}}_{1}{{\cal V}}_{2}}|\equiv\sqrt{\sum_{a}(e^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}})^{2}}$ (2)
are assigned for each oriented 1-simplex $a_{{{\cal V}}_{1}}a_{{{\cal V}}_{2}}$. The boundedness of the tetrad according to (2) is necessary for the convergence of the functional integral of the partition function. By this limitation imposed on the tetrad, the present work differs from the author’s previous works on the theory of diskette gravity. The conjecture is that the set of variables $\{\Omega,\,\hat{e}\}$ is an independent set of dynamic variables. Fermionic degrees of freedom (Dirac spinors) are assigned to each vertex of the complex:
$\Psi^{{\dagger}}_{{\cal V}},\quad\Psi_{{\cal V}}.$ (3)
The set of variables $\{\Psi^{{\dagger}},\,\Psi\}$ is a set of mutually independent variables, and the spinors $\Psi^{{\dagger}}_{{\cal V}}$ and $\Psi_{{\cal V}}$ are in mutual involution (or anti-involution) relative to Hermitian conjugation operation.
Consider a model with an action
$\mathop{{\mathfrak{A}}}\nolimits=\mathop{{\mathfrak{A}}}\nolimits_{g}+\mathop{{\mathfrak{A}}}\nolimits_{\Psi}+\mathop{{\mathfrak{A}}}\nolimits_{\Lambda_{0}}.$ (4)
Here $\mathop{{\mathfrak{A}}}\nolimits_{g}$ and $\mathop{{\mathfrak{A}}}\nolimits_{\Psi}$ are the actions of pure gravity and Dirac field, correspondingly:
$\mathop{{\mathfrak{A}}}\nolimits_{g}=-\frac{1}{5!\cdot 2\cdot l_{P}^{2}}\sum_{{\cal W}}\sum_{\sigma}\varepsilon_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})3})\sigma({{\cal V}}_{({{\cal W}})4})\sigma({{\cal V}}_{({{\cal W}})5})}$
×$\times\mathop{\rm tr}\nolimits\gamma^{5}\bigg{\{}\Omega_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}\Omega_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})}\Omega_{\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})5})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}\bigg{\}}.$ (5)
Each $\sigma$ is one of 5! vertex permutations ${{\cal V}}_{({{\cal W}})i}\longrightarrow\sigma({{\cal V}}_{({{\cal W}})i})$.
$\mathop{{\mathfrak{A}}}\nolimits_{\Psi}=\frac{1}{5\cdot 24^{2}}\sum_{{\cal W}}\sum_{\sigma}\varepsilon_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})3})\sigma({{\cal V}}_{({{\cal W}})4})\sigma({{\cal V}}_{({{\cal W}})5})}$
×$\times\mathop{\rm tr}\nolimits\gamma^{5}\bigg{\{}\hat{\Theta}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})2})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}\hat{e}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}\bigg{\}},$ (6)
$\hat{\Theta}_{{{\cal V}}_{1}{{\cal V}}_{2}}\equiv\Theta^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\gamma^{a}=\hat{\Theta}_{{{\cal V}}_{1}{{\cal V}}_{2}}^{{\dagger}},\quad\Theta^{a}_{{\cal V}_{1}{\cal V}_{2}}=\frac{i}{2}\left(\Psi^{{\dagger}}_{{\cal V}_{1}}\gamma^{a}\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\Psi_{{\cal V}_{2}}-\Psi^{{\dagger}}_{{\cal V}_{2}}\Omega_{{{\cal V}}_{2}{{\cal V}}_{1}}\gamma^{a}\Psi_{{\cal V}_{1}}\right).$ (7)
It is easy to check that (compare with (2))
$\hat{\Theta}_{{{\cal V}}_{1}{{\cal V}}_{2}}\equiv-\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\hat{\Theta}_{{{\cal V}}_{2}{{\cal V}}_{1}}\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}^{-1}.$ (8)
The contribution to the lattice action from the cosmological constant has the form
$\mathop{{\mathfrak{A}}}\nolimits_{\Lambda_{0}}=-\frac{1}{5!\cdot 12}\cdot\frac{\Lambda_{0}}{l_{P}^{2}}\varepsilon_{abcd}\sum_{{\cal W}}\sum_{\sigma}\varepsilon_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})3})\sigma({{\cal V}}_{({{\cal W}})4})\sigma({{\cal V}}_{({{\cal W}})5})}$
×$\times e^{a}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})1})}e^{b}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})2})}e^{c}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})3})}e^{d}_{\sigma({{\cal V}}_{({{\cal W}})5})\sigma({{\cal V}}_{({{\cal W}})4})}.$ (9)
The partition function is represented by the integral
$Z=\prod_{1-\mbox{simplices}}\int_{|e_{{{\cal V}}_{1}{{\cal V}}_{2}}|<1}\prod_{a}\mathop{\rm d}\nolimits e^{a}_{{\cal V}_{1}{\cal V}_{2}}\int\mathop{\rm d}\nolimits\mu\{\Omega_{{\cal V}_{1}{\cal V}_{2}}\}\prod_{{\cal V}}\int\mathop{\rm d}\nolimits\Psi^{{\dagger}}_{{\cal V}}\mathop{\rm d}\nolimits\Psi_{{\cal V}}\mathop{\rm exp}\nolimits(\mathop{{\mathfrak{A}}}\nolimits).$ (10)
Everywhere $\mathop{\rm d}\nolimits\mu\{\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\}$ is an invariant measure on the group $\mathop{\rm Spin}\nolimits(4)$.
The action (4) as well as integral (10) are invariant relative to the gauge transformations
$\tilde{\Omega}_{{{\cal V}}_{1}{{\cal V}}_{2}}=S_{{{\cal V}}_{1}}\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}S^{-1}_{{{\cal V}}_{2}},\quad\tilde{\hat{e}}_{{{\cal V}}_{1}{{\cal V}}_{2}}=S_{{{\cal V}}_{1}}\,\hat{e}_{{{\cal V}}_{1}{{\cal V}}_{2}}\,S^{-1}_{{{\cal V}}_{1}},\quad\tilde{\Psi}_{{\cal V}}=S_{{\cal V}}\Psi_{{\cal V}},\quad\tilde{\Psi^{{\dagger}}}_{{\cal V}}=\Psi_{{\cal V}}^{{\dagger}}S_{{\cal V}}^{-1},\quad S_{{\cal V}}\in\mathop{\rm Spin}\nolimits(4).$ (11)
Verification of this fact is facilitated by using the relation (compare with the relation for $\hat{e}_{{{\cal V}}_{1}{{\cal V}}_{2}}$ in (11))
$\tilde{\hat{\Theta}}_{{{\cal V}}_{1}{{\cal V}}_{2}}=S_{{{\cal V}}_{1}}\hat{\Theta}_{{{\cal V}}_{1}{{\cal V}}_{2}}S^{-1}_{{{\cal V}}_{1}},$ (12)
which follows directly from (11).
The considered lattice model is invariant with respect to the global discrete $\mathbb{Z}_{2}$ symmetry, which is an analog of the combined PT-symmetry. Let $\hat{\cal U}_{PT}$ denote the operator of this transformation. Then the transformed dynamic variables are expressed in terms of the original variables as follows:
$\hat{\cal U}_{PT}^{-1}\Psi_{{\cal V}}\hat{\cal U}_{PT}=U_{PT}\left(\Psi^{{\dagger}}_{{\cal V}}\right)^{t},\quad\hat{\cal U}_{PT}^{-1}\Psi^{{{\dagger}}}_{{\cal V}}\hat{\cal U}_{PT}=-\left(\Psi_{{\cal V}}\right)^{t}U^{-1}_{PT},\quad U_{PT}=i\gamma^{1}\gamma^{3}$
$\hat{\cal U}_{PT}^{-1}e^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\hat{\cal U}_{PT}=-e^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}},\quad\hat{\cal U}_{PT}^{-1}\omega^{ab}_{{{\cal V}}_{1}{{\cal V}}_{2}}\hat{\cal U}_{PT}=\omega^{ab}_{{{\cal V}}_{1}{{\cal V}}_{2}}.$ (13)
Here the superscript "$t"$ denotes the matrix transposition of the Dirac matrices and spinors. We have:
$U^{-1}_{PT}\gamma^{a}U_{PT}=(\gamma^{a})^{t},\quad U^{-1}_{PT}\sigma^{ab}U_{PT}=-(\sigma^{ab})^{t}.$ (14)
It follows from (13) and (14) that
$U^{-1}_{PT}\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}U_{PT}=\left(\Omega_{{{\cal V}}_{2}{{\cal V}}_{1}}\right)^{t},$ (15)
as well as
$\left(\Theta^{a}_{{\cal V}_{1}{\cal V}_{2}}\right)^{PT}=-\Theta^{a}_{{\cal V}_{1}{\cal V}_{2}}.$ (16)
II.2 The long-wavelength limit
Now let us pass on to the limit of slowly varying fields, when the action (4) reduces to the well known continuous gravity action. This transition have meaning together with the transition to Minkowski signature. As a result the compact gauge group $\mathop{\rm Spin}\nolimits(4)$ transforms into the non-compact group $\mathop{\rm Spin}\nolimits(3,1)$.
In this section, all lattice quantities for the Euclidean signature are primed. For the quantities related to the Minkowski signature, the previous designations are retained. This is done to make the transition from the Euclidean signature to the Minkowski signature transparent.
Firstly let us perform the following deformations of integration contours in integral (10):
${\omega^{\prime}}_{{{\cal V}}_{1}{{\cal V}}_{2}}^{4\alpha}=i\omega^{0\alpha}_{{{\cal V}}_{1}{{\cal V}}_{2}},\quad{\omega^{\prime}}_{{{\cal V}}_{1}{{\cal V}}_{2}}^{\alpha\beta}=-\omega_{{{\cal V}}_{1}{{\cal V}}_{2}}^{\alpha\beta},$
${e^{\prime}}^{4}_{{{\cal V}}_{1}{{\cal V}}_{2}}=e^{0}_{{{\cal V}}_{1}{{\cal V}}_{2}},\quad{e^{\prime}}^{\alpha}_{{{\cal V}}_{1}{{\cal V}}_{2}}=ie^{\alpha}_{{{\cal V}}_{1}{{\cal V}}_{2}}.$ (17)
The variables $\omega^{ab}_{{{\cal W}}ij}$, $e^{a}_{{{\cal W}}ij}$ are real quantities for Minkowski signature, and the indices take on the values
$a,\,b\ldots=0,1,2,3,\quad\alpha,\,\beta,\ldots=1,2,3.$ (18)
The Dirac matrices are transformed as follows:
${\gamma^{\prime}}^{4}=\gamma^{0},\quad{\gamma^{\prime}}^{\alpha}=i\gamma^{\alpha},\quad{\gamma^{\prime}}^{5}=\gamma^{5}=i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3},$
$\frac{1}{2}(\gamma^{a}\gamma^{b}+\gamma^{b}\gamma^{a})=\eta^{ab}=\mathop{\rm diag}\nolimits(1,\,-1,\,-1,\,-1),\quad\mathop{\rm tr}\nolimits\gamma^{5}\gamma_{a}\gamma_{b}\gamma_{c}\gamma_{d}=4i\varepsilon_{abcd},\quad\varepsilon_{0123}=1.$ (19)
Thus, for $\sigma^{ab}=(1/4)[\gamma^{a},\,\gamma^{b}]$ we get
${\sigma^{\prime}}^{4\alpha}=i\sigma^{0\alpha},\quad{\sigma^{\prime}}^{\alpha\beta}=-\sigma^{\alpha\beta}.$ (20)
Raising and lowering indices $a,b,\ldots$ is done using tensors $\eta^{ab}$ and $\eta_{ab}$, respectively. As a result of (17)-(20) we have
${\omega^{\prime}}_{{{\cal V}}_{1}{{\cal V}}_{2}}=\frac{1}{2}\omega^{ab}_{{{\cal V}}_{1}{{\cal V}}_{2}}\,\sigma_{ab}\equiv\omega_{{{\cal V}}_{1}{{\cal V}}_{2}},\quad{\hat{e}^{\prime}}_{{{\cal V}}_{1}{{\cal V}}_{2}}=\gamma_{a}e^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\equiv\hat{e}_{{{\cal V}}_{1}{{\cal V}}_{2}},$ (21)
and also
${\Omega^{\prime}}_{{{\cal V}}_{1}{{\cal V}}_{2}}=\mathop{\rm exp}\nolimits\left(\frac{1}{2}{\omega^{\prime}}_{{{\cal V}}_{1}{{\cal V}}_{2}}^{ab}{\sigma^{\prime}}^{ab}\right)=\mathop{\rm exp}\nolimits\left(\frac{1}{2}\omega_{{{\cal V}}_{1}{{\cal V}}_{2}}^{ab}\sigma_{ab}\right)\equiv\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\in\mathop{\rm Spin}\nolimits(3,1).$ (22)
We see that the holonomy elements $\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}$ become the elements of the non-compact group $\mathop{\rm Spin}\nolimits(3,1)$.
Dirac variables are transformed according to
$\Psi^{\prime}_{{\cal V}}=\Psi_{{\cal V}},\quad{\Psi^{\prime}}^{{\dagger}}_{{\cal V}}=\Psi_{{\cal V}}^{{\dagger}}\gamma^{0}=\overline{\Psi}_{{\cal V}}.$ (23)
in passing to the Minkowski signature.
Now let us pass on to the formal limit of slowly varying fields, when the action (4) reduces to the well known continuous gravity action. In order to pass to the long-wavelength limit let’s consider a certain 4D sub-complex ${\mathop{{\mathfrak{K}}}\nolimits}^{\prime}\in\mathop{{\mathfrak{K}}}\nolimits$ with the trivial topology of four-dimensional disk. Realize geometrically this sub-complex in $\mathbb{R}^{4}$. Suppose that the geometric realization is the triangulation of a compact part of $\mathbb{R}^{4}$, so that in $\mathbb{R}^{4}$ we have $(\mbox{int}\,s^{4}_{{\cal W}})\cap(\mbox{int}\,s^{4}_{{{\cal W}}^{\prime}})=\emptyset$ for ${{\cal W}}\neq{{\cal W}}^{\prime}$ and the sizes of these simplices are commensurable. Thus each vertex of the sub-complex acquires the coordinates $x^{\mu}$ which are the coordinates of the vertex image in $\mathbb{R}^{4}$:
[…]
We stress that the correspondence between the vertices from the considered subset and the coordinates (24) is one-to-one.
Consider a certain simplex $s^{4}_{{\cal W}}$. We denote all five vertices of this 4-simplex as ${{\cal V}}_{i}$, $i=1,2,3,4$ and ${{\cal V}}_{m}\neq{{\cal V}}_{i}$. Note that each of these vertices also belongs to other adjacent 4-simplexes. From the aforesaid it is evident that the four infinitesimal vectors
[…]
are linearly independent. The vector system (25) refers to the simplex $s^{4}_{{\cal W}}$ and its vertex ${{\cal V}}_{m}$, while each of the vectors of the system (25) belongs to several adjacent 4-simplexes. The differentials of coordinates (25) correspond to one-dimensional simplices $a_{{{\cal V}}_{m}}a_{{{\cal V}}_{i}}$, so that, if the vertex $a_{{{\cal V}}_{m}}$ has coordinates $x^{\mu}_{{{\cal V}}_{m}}$, then the vertex $a_{{{\cal V}}_{i}}$ has the coordinates $x^{\mu}_{{{\cal V}}_{m}}+\mathop{\rm d}\nolimits x^{\mu}_{\left({{\cal V}}_{m}{{\cal V}}_{i}\right)}$.
In the continuous limit, the holonomy group elements $\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}$ are close to the identity element, so that the quantities $\omega^{ab}_{{{\cal V}}_{1}{{\cal V}}_{2}}$ tend to zero being of the order of $O(\mathop{\rm d}\nolimits x^{\mu})$ and they slowly change when moving along the lattice. Thus one can consider the following system of equation for $\omega_{\mu}\left({\cal W},{{\cal V}}_{m},{{\cal V}}_{i}\right)$:
[…]
In this system of linear equation, the indices ${{\cal W}}$ and $m$ are fixed, the summation is carried out over the index $\mu$. Since the vectors (25) are linearly independent, the quantities $\omega_{\mu}\left({\cal W},{{\cal V}}_{m},{{\cal V}}_{i}\right)$ are defined uniquely.
The problem is the definition of the connection 1-form field $\omega_{\mu}\big{(}(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})/2\big{)}$ satisfying systems of equations (26) for all 1-simplexes $a_{{{\cal V}}_{m}}a_{{{\cal V}}_{i}}$ independently of 4-simplexes and the vertices marked in them. The inverse problem is solved obviously: if a 1-form $\omega_{\mu}\big{(}(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})/2\big{)}$ is given, then the set of values $\omega_{{{\cal V}}_{m}{{\cal V}}_{i}}$ is determined according to (26).
Note that due to Eqs. (2), (25) and (26) we have
$\omega_{\mu}\left({\cal W},{{\cal V}}_{m},{{\cal V}}_{i}\right)\,\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\omega_{\mu}\left({\cal W},{{\cal V}}_{i},{{\cal V}}_{m}\right)\,\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}.$ (27)
Suppose that a one-dimensional simplex $a_{{{\cal V}}_{m}}a_{{{\cal V}}_{i}}$ belongs to four-dimensional simplices with indices ${{\cal W}}_{1},\,{{\cal W}}_{2},\,\ldots\,,\,{{\cal W}}_{r}$. Introduce the quantity
$\omega_{\mu}\left(\frac{1}{2}\,(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})\,\right)\equiv\frac{1}{2r}\sum_{s=1}^{r}\bigg{\{}\omega_{\mu}\left({{\cal W}}_{s},{{\cal V}}_{m},{{\cal V}}_{i}\right)+\omega_{\mu}\left({{\cal W}}_{s},{{\cal V}}_{i},{{\cal V}}_{m}\right)\bigg{\}}\,,$ (28)
which is assumed to be related to the midpoint of the segment $[x^{\mu}_{{{\cal V}}_{m}},\,x^{\mu}_{{{\cal V}}_{i}}\,]$. According to the definition, we have the following chain of identities:
$\omega_{{{\cal V}}_{({{\cal W}}_{1})m}{{\cal V}}_{({{\cal W}}_{1})i}}\equiv\ldots\equiv\omega_{{{\cal V}}_{({{\cal W}}_{r})m}{{\cal V}}_{({{\cal W}}_{1})r}}.$ (29)
It follows from (25)–(29) that
$\omega_{\mu}\left(\frac{1}{2}\,(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})\,\right)\,\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\omega_{{{\cal V}}_{m}{{\cal V}}_{i}}\,.$ (30)
The value of the field element $\omega_{\mu}$ in (30) is uniquely defined by the midpoint of corresponding one-dimensional simplex. At any points, the connection 1-form is determined by continuity.
Hence, we assume that the fields $\omega_{\mu}$ smoothly depend on the points belonging to the geometric realization of each four-dimensional simplex. In this case, the following formula is valid up to $O\big{(}(\mathop{\rm d}\nolimits x)^{2}\big{)}$ inclusive
$\Omega_{{{\cal V}}_{m}{{\cal V}}_{i}}\Omega_{{{\cal V}}_{i}{{\cal V}}_{j}}\Omega_{{{\cal V}}_{j}{{\cal V}}_{m}}=\mathop{\rm exp}\nolimits\left[\frac{1}{2}\,\mathop{{\mathfrak{R}}}\nolimits_{\mu\nu}(x_{{{\cal V}}_{m}})\mathop{\rm d}\nolimits x^{\mu}_{\left({{\cal V}}_{m}{{\cal V}}_{i}\right)}\mathop{\rm d}\nolimits x^{\nu}_{\left({{\cal V}}_{m}{{\cal V}}_{j}\right)}\right],$ (31)
where
$\mathop{{\mathfrak{R}}}\nolimits_{\mu\nu}=\partial_{\mu}\omega_{\nu}-\partial_{\nu}\omega_{\mu}+[\omega_{\mu},\,\omega_{\nu}\,]\,.$ (32)
In exact analogy with (26), let us write out the following relations for a tetrad field without explanations:
$\hat{e}_{\mu}\left({\cal W},{{\cal V}}_{m},{{\cal V}}_{i}\right)\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\hat{e}_{{{\cal V}}_{m}{{\cal V}}_{i}},\quad i=1,2,3,4.$ (33)
So 1-form
$e^{a}(x)=e^{a}_{\mu}(x)\mathop{\rm d}\nolimits x^{\mu}$ (34)
arises.
With the help of Equations given in this Section, we rewrite the lattice action (4) in the long-wavelength limit:
${\mathop{{\mathfrak{A}}}\nolimits^{\prime}}_{g\,Lat}\longrightarrow i\mathop{{\mathfrak{A}}}\nolimits_{g},\quad\mathop{{\mathfrak{A}}}\nolimits_{g}=-\frac{1}{4\,l^{2}_{P}}\varepsilon_{abcd}\int\mathop{{\mathfrak{R}}}\nolimits^{ab}\wedge e^{c}\wedge e^{d},\quad\mathop{{\mathfrak{R}}}\nolimits^{ab}=\mathop{{\mathfrak{R}}}\nolimits^{ab}_{\mu\nu}\mathop{\rm d}\nolimits x^{\mu}\wedge\mathop{\rm d}\nolimits x^{\nu},$ (35)
${\mathop{{\mathfrak{A}}}\nolimits^{\prime}}_{\Psi\,Lat}\longrightarrow i\mathop{{\mathfrak{A}}}\nolimits_{\Psi},\quad\mathop{{\mathfrak{A}}}\nolimits_{\Psi}=\frac{1}{6}\varepsilon_{abcd}\int\Theta^{a}\wedge e^{b}\wedge e^{c}\wedge e^{d},$
$\Theta^{a}=\frac{i}{2}\left[\overline{\Psi}\gamma^{a}{\cal D}_{\mu}\,\Psi-\left(\overline{{\cal D}_{\mu}\,\Psi}\right)\gamma^{a}\Psi\right]\mathop{\rm d}\nolimits x^{\mu},\quad{\cal D}_{\mu}=\left(\partial_{\mu}+\omega_{\mu}\right),$ (36)
$\mathop{{\mathfrak{A}}}\nolimits^{\prime}_{\Lambda_{0}\,Lat}\longrightarrow i\mathop{{\mathfrak{A}}}\nolimits_{\Lambda_{0}},\quad\mathop{{\mathfrak{A}}}\nolimits_{\Lambda_{0}}=-\frac{2\Lambda_{0}}{l_{P}^{2}}\int e^{0}\wedge e^{1}\wedge e^{2}\wedge e^{3}.$ (37)
In formulas (35)-(37) the additional subscript "$Lat"$ means that the corresponding values are defined on the lattice.
In the long-wave limit, there is a general covariance of the theory: all terms of the action (35)-(37) are invariant relative to general coordinate transformations.
In the Minkowski signature, the PT-symmetry of the action is determined by the formulas (13)-(14), (16) with the only difference that in these formulas the replacement should be made $\Psi^{{\dagger}}\longrightarrow\overline{\Psi}$.
III Superhigh temperatures in the lattice model of gravity
Now let’s simulate the superhigh temperature in the lattice theory. As is known, in quantum theory, the partition function $\mathop{\rm tr}\nolimits\mathop{\rm exp}\nolimits(-\beta{\cal H})$ differs from the transition amplitude over a finite time $\Delta t$, represented as a functional integral, by the transition to imaginary time $\Delta t=-i\beta$, $\beta=1/T$, and taking the trace. We must study some properties of the partition function (10) in the case of ultra-high temperatures, that is, $\beta\longrightarrow 0$.
When applied to the lattice theory studied here, this means the following.
Suppose that a D-dimensional lattice has two (D-1)-dimensional sublattices $\Sigma_{1}$ and $\Sigma_{2}$ that form its boundary. For simplicity, we assume that between $\Sigma_{1}$ and $\Sigma_{2}$ there are $N<\infty$ 4D lattice layers, each one edge thick. Between $\Sigma_{1}$ and $\Sigma_{2}$ there is a 3D sublattice $\Sigma_{0}$, and the entire 4D lattice is symmetrical with respect to the mirror reflection relative to the $\Sigma_{0}$ sublattice. It follows that the sublattices $\Sigma_{1}$ and $\Sigma_{2}$ are identical, which makes it possible to calculate the partition function.
Let $\Phi_{1\xi}$ and $\Phi_{2\xi}$ denote the holomorphic functions of fermion variables with real coefficients defined on $\Sigma_{1}$ and $\Sigma_{2}$, respectively. To calculate the trace in fermionic variables, it is necessary to use the full set of holomorphic wave functions $\Phi_{1\xi}\equiv\Phi_{1\xi}\{\Psi^{{\dagger}}_{{{\cal V}}}\}$, and their Hermitian conjugate functions $\Phi_{2\xi}^{{\dagger}}\equiv\left(\Phi_{2\xi}\{\Psi^{{\dagger}}_{{{\cal V}}}\}\right)^{{\dagger}}=\Phi_{2\xi}\{\Psi_{{{\cal V}}}\}$. Here the index $\xi$ lists the independent wave functions from their complete set. The functional
$\Phi_{2\xi}^{{\dagger}}\mathop{\rm exp}\nolimits(\beta\mathop{{\mathfrak{A}}}\nolimits)\Phi_{1\xi}$ (38)
should be placed under the integral (10) and the sum over $\xi$ taken. In the integral (10) on the identified 1-simplices belonging to $\Sigma_{1}$ and $\Sigma_{2}$, the variables $\{\Omega\}$ and $\{e^{a}\}$ must be identified .
We are interested in the case $\beta\ll 1$. We intend to prove that discrete PT symmetry (13)-(16) is not broken at ultrahigh temperatures.
The following Statement holds:
Statement. In a certain finite neighborhood of the point $\beta=0$, the free energy of the partition function (10), with the exception of the term of the form $(-(C_{1}\cdot\mathop{{\mathfrak{M}}}\nolimits+4\mathop{{\mathfrak{N}}}\nolimits\nu_{\Psi})\cdot\ln\beta)$ (see (46)), is a holomorphic function of the variable $\beta$. All action symmetries (4), including discrete PT symmetries, are preserved. $\square$
Let us give some arguments in favor of the Statement. Let us consider the high-temperature expansion of the partition function in the 2D Ising model, which is the sum over all closed paths with self-intersections on the lattice. Let the lattice contain $\mathop{{\mathfrak{L}}}\nolimits\longrightarrow\infty$ edges and let a path (possibly several non-intersecting paths) with a total of $\mathop{{\mathfrak{l}}}\nolimits$ edges be fixed on it. Note that $\mathop{{\mathfrak{l}}}\nolimits$ is an even number and $\mathop{{\mathfrak{l}}}\nolimits\geq 4$. When calculating the partition function, it is convenient to isolate the factor $(\cosh\beta)^{\mathop{{\mathfrak{L}}}\nolimits}$, which does not give a singularity in the free energy. Then each edge of the path should be assigned a weight $(\tanh\beta)$. Let $g_{\mathop{{\mathfrak{l}}}\nolimits}$ be the number of such closed paths. We have a strict inequality
$g_{\mathop{{\mathfrak{l}}}\nolimits}<\frac{\mathop{{\mathfrak{L}}}\nolimits!}{\mathop{{\mathfrak{l}}}\nolimits!(\mathop{{\mathfrak{L}}}\nolimits-\mathop{{\mathfrak{l}}}\nolimits)!},$ (39)
since the number on the right side of the inequality (39) also includes the number of all non-closed paths of weight $(\tanh\beta)^{\mathop{{\mathfrak{l}}}\nolimits}$. Therefore, the Ising partition function $Z$
$Z/(\cosh\beta)^{\mathop{{\mathfrak{L}}}\nolimits}=1+\sum_{\mathop{{\mathfrak{l}}}\nolimits}g_{\mathop{{\mathfrak{l}}}\nolimits}(\tanh\beta)^{\mathop{{\mathfrak{l}}}\nolimits}<\sum_{\mathop{{\mathfrak{l}}}\nolimits=0}^{\mathop{{\mathfrak{L}}}\nolimits}\frac{\mathop{{\mathfrak{L}}}\nolimits!}{\mathop{{\mathfrak{l}}}\nolimits!(\mathop{{\mathfrak{L}}}\nolimits-\mathop{{\mathfrak{l}}}\nolimits)!}(\tanh\beta)^{\mathop{{\mathfrak{l}}}\nolimits}$
$=(1+\tanh\beta)^{\mathop{{\mathfrak{L}}}\nolimits}\equiv Z_{M}.$ (40)
This shows that the specific free energy per edge $f_{M}=-\ln Z_{M}/\mathop{{\mathfrak{L}}}\nolimits=-\ln(1+\tanh\beta)$ is a holomorphic function in a finite neighborhood of the point $\beta=0$ and in the case of $\mathop{{\mathfrak{L}}}\nolimits\longrightarrow\infty$. This observation is a consequence of the fact that the quantity $Z_{M}$ is essentially local for small $\beta$, that is, effectively $Z_{M}$ is the product of local holomorphic functions. In other words, there is no long-range order at small $\beta$. But if open contours are eliminated from the sum $Z_{M}$, then long-range order will not arise. However, the size of the holomorphy neighborhood may change. This means that the limit $f=-\ln Z/\mathop{{\mathfrak{L}}}\nolimits$ at $\mathop{{\mathfrak{L}}}\nolimits\longrightarrow\infty$ exists. Indeed, the value on the left side (40) can be represented as
$Z/(\cosh\beta)^{\mathop{{\mathfrak{L}}}\nolimits}=[1+\rho(\tanh\beta)]^{\mathop{{\mathfrak{L}}}\nolimits},$ (41)
and the holomorphic function $\rho(\tanh\beta)\rightarrow 0$ for $\beta\rightarrow 0$. That’s why
$f=-\ln\cosh\beta-\ln[1+\rho(\tanh\beta)]$ (42)
is a holomorphic function in some finite neighborhood of the point $\beta=0$.
Similar reasoning can be carried out in the case of a partition function (10), although this case is qualitatively more complicated. Both the fermion and bosonic integrals in (10) are well defined. Let us consider qualitatively the results of these integrations.
Let ${\mathop{{\mathfrak{N}}}\nolimits}\longrightarrow\infty$ and $\nu_{\Psi}$ denote the numbers of vertices of the complex and Dirac fields, respectively. Integration into (10) over the Dirac field gives the factor under the remaining integrals of the form
$\beta^{4{\mathop{{\mathfrak{N}}}\nolimits}\nu_{\Psi}}{\mathop{{\mathfrak{P}}}\nolimits}\{\Omega,e^{a}\},$ (43)
where ${\mathop{{\mathfrak{P}}}\nolimits}\{\Omega,e^{a}\}$ is a gauge-invariant homogeneous polynomial of the variables $\Omega$ and $e^{a}$ of degrees ${4{\mathop{{\mathfrak{N}}}\nolimits}\nu_{\Psi}}$ and ${12{\mathop{{\mathfrak{N}}}\nolimits}\nu_{\Psi}}$, respectively.
Next, we consider integrals over bosonic variables. First we consider integration over the connection variables $\Omega$.
Let the polynomial ${\mathop{{\mathfrak{P}}}\nolimits}\{\Omega,e^{a}\}$ contain a variable $\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}$ to an odd power. Integral $(\int\mathop{\rm d}\nolimits\mu\{\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\}\ldots)$ of the tensor product of an odd number of elements $\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}$ is equal to zero. Let us expand the exponential into terms containing the element $\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}$. Let us assume that as a result of a finite number of such steps (for a finite subcomplex) we will have a tensor product of an even number of elements $\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}$ on each 1-simplex $a_{{{\cal V}}_{1}}a_{{{\cal V}}_{2}}$. Therefore, as a result of integration over the gauge group, the first nonzero term will receive an additional factor $\beta^{C\cdot\mathop{{\mathfrak{M}}}\nolimits}$, where $\mathop{{\mathfrak{M}}}\nolimits\rightarrow\infty$ is the number of 1-simplices of the complex and $0\leq C\sim 1$. All other terms of this expansion, as a result of integration over the gauge group, give an additional gauge invariant term $\mathop{{\mathfrak{F}}}\nolimits\{e^{a};\beta\}$ in the partition function. This term is a functional of the variables $\{e^{a}\}$ and a well-converging holomorphic function of the parameter $\beta$ in the neighborhood of zero, and
$\frac{\mathop{{\mathfrak{F}}}\nolimits\{e^{a};\beta\}}{\beta^{C\cdot\mathop{{\mathfrak{M}}}\nolimits+4\mathop{{\mathfrak{N}}}\nolimits\nu_{\Psi}}}\rightarrow 0\quad\mbox{as}\quad\beta\rightarrow 0.$
There are only two types of invariants relative to gauge transformations (11) associated with a vertex $a_{{\cal V}}$:
$e^{a}_{{{\cal V}}{{\cal V}_{1}}}e^{a}_{{{\cal V}}{{\cal V}_{2}}},$ (44)
$\varepsilon_{abcd}e^{a}_{{{\cal V}}{{\cal V}_{1}}}e^{b}_{{{\cal V}}{{\cal V}_{2}}}e^{c}_{{{\cal V}}{{\cal V}_{3}}}e^{d}_{{{\cal V}}{{\cal V}_{4}}}.$ (45)
In (44) the vertices $a_{{{\cal V}}_{1}}$ and $a_{{{\cal V}}_{2}}$ are not necessarily different, and in (45) the 1-simplices $(a_{{\cal V}}a_{{{\cal V}}_{1}})$, $(a_{{\cal V}}a_{{{\cal V}}_{2}})$, $(a_{{\cal V}}a_{{{\cal V}}_{3}})$, $(a_{{\cal V}}a_{{{\cal V}}_{4}})$ do not necessarily belong to the same 4-simplex.
Let us briefly consider the integral over variables $\{e^{a}\}$. Note that if under the integral there is an expression (44) to the first degree and it does not intersect with other similar expressions, then the integral is identically equal to zero. A similar statement holds for the expression (45). Therefore, under the integral the expressions (44) and (45) must be in such powers and combinations that the variables $e^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}$ on each 1-simplex are in an even power (including zero). For this purpose, expansion of the exponential in (38) in the contribution to the action $\beta\mathop{{\mathfrak{A}}}\nolimits_{\Lambda_{0}}$ (see (9)) may be necessary.
The action (4) is local, that is, it consists of terms (local operators), each of which is defined on the nearest lattice elements. Let’s call one of these operators $A$, and the other $B$, and we will assume that the operators $A$ and $B$ are separated by a considerable distance along the lattice. Consider the correlator $\langle AB\rangle$. For this correlator to be non-zero, it is necessary to expand the exponential in these operators so that between the operators $A$ and $B$ there is a sufficient number of adjacent local operators. But such an expansion will give the parameter $\beta\ll 1$ to a fairly high degree, at least proportional to the number of 1-simplices between $A$ and $B$, and with a coefficient of the order of unity. Therefore, the partition function in the high-temperature phase is local in nature.
Let us draw a conclusion from the above consideration. At ultrahigh temperature $(\beta=1/T\longrightarrow 0)$ the integral for the partition function (10) is a holomorphic function in a neighborhood of the point $\beta=0$ and has the form (compare with (41))
$Z=\mathop{\rm Const}\nolimits\cdot\beta^{C_{1}\cdot\mathop{{\mathfrak{M}}}\nolimits+4\mathop{{\mathfrak{N}}}\nolimits\nu_{\Psi}}(1+\mathop{{\mathfrak{f}}}\nolimits(\beta))^{C_{2}\cdot\mathop{{\mathfrak{M}}}\nolimits},$
$\mathop{{\mathfrak{f}}}\nolimits(\beta)\rightarrow 0\quad\mbox{as}\quad\beta\rightarrow 0,\quad C_{1}\sim C_{2}\sim 1.$ (46)
This implies the validity of the Statement. In turn, it follows from the Statement that in the high-temperature phase no symmetries are broken, including discrete PT symmetry.
Based on the validity of the Statement, we will begin to calculate the averages of some operators.
The discrete PT transformation (13)-(16) does not change the measure and integrand (except, perhaps, the averaged value), but swaps the initial and final states:
$\left(\Phi_{2\xi}^{{\dagger}}\right)^{\prime}\equiv\Phi_{2\xi}^{{\dagger}}\hat{\cal U}_{PT}^{{\dagger}}=\Phi_{2\xi}\{U_{PT}(\Psi_{{\cal V}}^{{\dagger}})^{t}\},$
$\Phi_{1\xi}^{\prime}\equiv\hat{\cal U}_{PT}\Phi_{1\xi}=\Phi_{1\xi}\{-\Psi^{t}_{{\cal V}}U^{-1}_{PT}\},$
$\left(\Phi_{2\xi}^{{\dagger}}\right)^{\prime}\cdot\Phi_{1\xi}^{\prime}=\Phi_{1\xi}\{\Psi^{t}_{{\cal V}}U^{-1}_{PT}\}\cdot\Phi_{2\xi}\{U_{PT}(\Psi_{{\cal V}}^{{\dagger}})^{t}\}.$ (47)
Here, a prime above the wave function symbol means that it is PT-transformed. From the last equality it is clear that the initial and final wave functions are swapped as a result of the PT transformation, and their scalar product is preserved. This means that the operator $\hat{\cal U}_{PT}$ is anti-unitary.
We have a chain of equalities:
$\langle{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\rangle_{e,\Psi}=N\sum_{\xi}\int_{e}\int_{\Psi^{{\dagger}}\Psi}\Phi_{2\xi}^{{\dagger}}{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\mathop{\rm exp}\nolimits(\beta\mathop{{\mathfrak{A}}}\nolimits)\Phi_{1\xi}$
$=N\sum_{\xi}\int_{e}\int_{\Psi^{{\dagger}}\Psi}\left(\Phi_{2\xi}^{{\dagger}}\right)^{\prime}{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\mathop{\rm exp}\nolimits(\beta\mathop{{\mathfrak{A}}}\nolimits)\Phi_{1\xi}^{\prime}$
$=N\sum_{\xi}\int_{e}\int_{\Psi^{{\dagger}}\Psi}\Phi_{1\xi}^{{\dagger}}\left\{\hat{\cal U}_{PT}^{-1}{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\mathop{\rm exp}\nolimits(\beta\mathop{{\mathfrak{A}}}\nolimits)\hat{\cal U}_{PT}\right\}^{{\dagger}}\Phi_{2\xi}$
$=-N\sum_{\xi}\int_{e}\int_{\Psi^{{\dagger}}\Psi}\Phi_{1\xi}^{{\dagger}}{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\mathop{\rm exp}\nolimits(\beta\mathop{{\mathfrak{A}}}\nolimits)\Phi_{2\xi}=-\langle{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\rangle_{e,\Psi}.$ (48)
Here $N$ is a normalizing constant, the integral is calculated only over the tetrad and Dirac field variables, but not over the connection variables. If the integration also included an integral over the variables $\{\Omega\}$, then any gauge non-invariant quantity, like ${\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}$, would vanish identically. But the integral (48) is meaningful, since the gauge is fixed in it. In (48), the first equality is the definition of the average value, the second equality is a consequence of the Statement, the third and fourth equalities are consequences of the equalities (47) and (16), respectively. The last equality follows from the mirror symmetry of the complex with respect to the 3D subcomplex $\Sigma_{0}$.
The main conclusion of this work follows from the chain of equalities (48):
$\langle{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\rangle_{e,\Psi}\equiv\langle{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\rangle_{\mbox{Gauge Fix}}=0.$ (49)
The calculation of the vacuum mean value ${\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}$ in the continuum limit in the Minkowski signature at zero temperature was carried out in the work Vergeles (2021b):
$\langle 0|\Theta^{a}_{\mu}|0\rangle=\frac{2}{\pi^{2}l^{4}_{P}}e^{a}_{\mu}\neq 0,$
$\Theta^{a}_{\mu}=\frac{i}{2}\left[\overline{\Psi}\gamma^{a}{\cal D}_{\mu}\,\Psi-\left(\overline{{\cal D}_{\mu}\,\Psi}\right)\gamma^{a}\Psi\right],\quad{\cal D}_{\mu}=\left(\partial_{\mu}+\omega_{\mu}\right).$ (50)
Comparison of Eqs (49) and (50) shows that in the lattice theory of gravity coupled with Dirac field there is a phase transition in temperature.
Phase transition from the zero value of the bilinear form (with respect to Dirac fields) $\Theta^{a}_{\mu}$ to its nonzero value and the meaning of this transition for the 4D theory of gravity are discussed in Volovik (2022). In this regard, an article Volovik (1990) may also be of interest.
IV Einstein equation and solution
Here the standard spatially flat Robertson-Walker metric
$\mathop{\rm d}\nolimits s^{2}=\mathop{\rm d}\nolimits t^{2}-a^{2}(t)(\mathop{\rm d}\nolimits x^{\alpha})^{2},\quad\alpha=1,2,3$ (51)
is used, $H={\dot{a}}/a$ is Hubble constant and $a(t)$ is the cosmic scale factor. We assume $c=1$. Appropriate remarks are made in formulas with restored dimension.
First of all, we will show that at the end of the inflation process and further, vacuum quantum fluctuations are insignificant when describing the dynamics of macroscopic regions of space and the matter contained in them.
Our approach assumes that all physical quantities are determined by taking into account vacuum quantum zero point fluctuations. In particular, the energy density and pressure are mainly determined by quantum fluctuations. It is well known Schwinger (2018) that the following simultaneous commutation relations for the components of the energy-momentum tensor take place in the Minkowski space:
$\left[T^{00}({\bf x}),T^{00}({\bf y})\right]=-i\hbar\left(T^{0k}({\bf x})+T^{0k}({\bf y})\right)\partial_{k}\delta^{(3)}({\bf x}-{\bf y})+\mbox{Schwinger terms},$ (52)
and so on. The Schwinger terms are the higher derivatives of the $\delta$-functions and are not of interest here. The same commutation relations take place in a curved space if they are written near the center of Riemann’s normal coordinates.
Let us denote by $l_{P}$ the minimum possible size on the lattice, which we will call the Planck size. Then $\delta^{(3)}({\bf 0})\sim l_{P}^{-3}$. Thus we find, using (52) and the well-known rule, the order of quantum fluctuations of the energy density $\Delta\varepsilon$ at wavelengths $\lambda\gg l_{P}$:
$\Delta\varepsilon\sim\sqrt{\frac{\hbar|\varepsilon|}{l_{P}^{3}\lambda}},\quad\frac{\Delta\varepsilon}{|\varepsilon|}\sim\sqrt{\frac{\hbar}{l_{P}^{3}\lambda|\varepsilon|}}.$ (53)
In the modern era and also at the end of the inflation phase, the vacuum temperature $T_{\mbox{vac}}\sim\hbar H$ (see Section V) is very small compared to the maximum (modulo) particle energy in the Dirac sea $|\epsilon|\sim\hbar/l_{P}$. It follows that the vacuum energy density $|\varepsilon|\sim\hbar/l_{P}^{4}$, and we obtain the estimation
$\Delta\varepsilon/|\varepsilon|\sim\sqrt{l_{P}/\lambda}\sim 10^{-16}\quad\mbox{for}\quad\lambda\sim 1\,\mbox{cm},\quad l_{P}\sim 10^{-32}\,\mbox{cm}.$ (54)
Note that, provided that the long-wave limit is valid, it makes sense to assume $\lambda/a\sim\mathop{\rm const}\nolimits_{1}$, $l_{P}/a\sim\mathop{\rm const}\nolimits_{2}$, and therefore $l_{P}/\lambda\sim\mathop{\rm const}\nolimits$. Therefore the assessment (54) remains valid over a very wide range, including a significant part of the inflation phase. This shows that the use of Einstein’s classical equations to describe macroscopic dynamics in the modern era is justified. However, when approaching the Big Bang point, the vacuum temperature becomes very high, as a result of which the energy density $|\varepsilon|$ quickly decreases in absolute value due to the population of positive energy levels by fermions. Estimates (53) and (54) become unsatisfactory, and in the situation $\Delta\varepsilon/|\varepsilon|\sim 1$ classical equations are unsuitable, the dynamics become purely quantum. A detailed study of the boundary of the transition from classical to quantum dynamics is not included in the scope of this work.
Let us demonstrate the tendency of the average energy of the system to zero with increasing temperature using a simple example. Consider one Fermi particle, which can be in only two states with energies ±$\pm\epsilon$. Suppose that the particle is placed in a thermostat, so its chemical potential is zero and the temperature $T$ tends to infinity ($\epsilon/T\longrightarrow 0$). Then $\varepsilon=<\epsilon>\longrightarrow 0$.
IV.1 Solution of Einstein’s equations with bare cosmological constant
We assume that the energy-momentum tensor of an ideal relativistic fluid is suitable for solving the problem posed:
$T^{a}_{b}=(\varepsilon+p)U^{a}U_{b}-p\delta^{a}_{b}.$ (55)
We work in an orthonormal basis in which the metric tensor $\eta_{ab}=\mathop{\rm diag}\nolimits(1,-1,-1,-1)$. On the right side of (55), the symbols $\varepsilon$ and $p$ denote the energy density and pressure, respectively, and these quantities also include vacuum energy and pressure. Since fermion fields, in contrast to bosonic ones, make a negative contribution to the vacuum energy, but there are significantly more fermionic degrees of freedom than bosonic ones, we have $\varepsilon<0$. Moreover, lattice regularization means that $|\varepsilon|,|p|<\infty$. $U^{a}$ is the averaged 4-velocity of the macroscopic regions of the lattice. In our case $U^{a}=(1,0,0,0)$. To compensate for the vacuum energy, a bare finite positive cosmological constant $\Lambda_{0}$ is introduced into the Einstein equation:
$\mathop{{\mathfrak{R}}}\nolimits^{a}_{b}-\frac{1}{2}\delta^{a}_{b}\mathop{{\mathfrak{R}}}\nolimits=8\pi G\,T^{a}_{b}+\Lambda_{0}\delta^{a}_{b}.$ (56)
In lattice theory, the cosmological constant is introduced in a natural way (see (9)). We assume that the bare cosmological constant (with restored dimension)
$\Lambda_{0}=\mathop{\rm const}\nolimits\sim l_{P}^{-2},\quad l_{P}\sim\sqrt{\frac{8\pi G\hbar}{c^{3}}}\sim 10^{-32}\mbox{cm}.$ (57)
Note that the bare cosmological constant on the lattice (9) is dimensionless, like all other constants and variables, and it is of the order of unity.
For the metric, we use ansatz (51). In order not to clutter up the formulas, we introduce the notation
$8\pi G\,\varepsilon=\tilde{\varepsilon},\quad 8\pi G\,p=\tilde{p}.$ (58)
All components of the Einstein equation are reduced to two equations:
$3\frac{\dot{a}^{2}}{a^{2}}=\Lambda_{0}+\tilde{\varepsilon},\quad 2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}=\Lambda_{0}-\tilde{p}.$ (59)
Another equation $\nabla_{a}T^{a}_{b}=0$ is a consequence of equations (59), and therefore it does not need to be considered. With the help of the Hubble constant $H(t)$ Eqs. (59) are rewritten as follows:
$0\displaystyle 2\dot{H}+(\tilde{\varepsilon}+\tilde{p})=0,\quad 3H^{2}-(\Lambda_{0}+\tilde{\varepsilon})=$ 0 . (60)
So we have 3 unknown functions $\{\tilde{\varepsilon}(t),\tilde{p}(t),H(t)\}$ and 2 equations (60). The missing equation is the equation of state relating energy density and pressure. Regarding the equation of state, the following facts are reliably known: (i) in the case of an ideal Boltzmann gas, we have $\tilde{p}=(2/3)\tilde{\varepsilon}$; (ii) in the case of real dusty matter, we have $\tilde{p}=0$; (iii) in the case of real ultrarelativistic matter we have $\tilde{p}=\tilde{\varepsilon}/3$; (iiii) in the case of vacuum energy and pressure, we have $\tilde{p}=-\tilde{\varepsilon}$. In all four cases, the energy density and pressure are linearly related. In addition, these quantities have the same dimension. Therefore, we propose to accept the following hypothesis:
$\tilde{p}=\varkappa\Lambda_{0}+(\varkappa-1)\tilde{\varepsilon}\longleftrightarrow\tilde{\varepsilon}+\tilde{p}=\varkappa(\tilde{\varepsilon}+\Lambda_{0}).$ (61)
This equation is linear and inhomogeneous with an unknown function $\varkappa(t)$, the asymptotics of which are further determined based on the known dynamics. The set of equations (60) and (61) has a solution:
$\dot{H}=-\frac{3}{2}\varkappa H^{2}\longrightarrow$
$H(t)=H_{0}\left(1+\frac{3}{2}H_{0}\int_{t_{0}}^{t}\varkappa(t^{\prime})\mathop{\rm d}\nolimits t^{\prime}\right)^{-1},$ (62)
$\tilde{\varepsilon}(t)=-\Lambda_{0}+3H_{0}^{2}\left(1+\frac{3}{2}H_{0}\int_{t_{0}}^{t}\varkappa(t^{\prime})\mathop{\rm d}\nolimits t^{\prime}\right)^{-2},$ (63)
$\tilde{p}(t)=\Lambda_{0}+3\big{(}\varkappa(t)-1\big{)}H_{0}^{2}\left(1+\frac{3}{2}H_{0}\int_{t_{0}}^{t}\varkappa(t^{\prime})\mathop{\rm d}\nolimits t^{\prime}\right)^{-2}.$ (64)
Here $H_{0}$ is the integration constant which plays the role of the Hubble constant at the beginning of the inflation phase. The moment $t_{0}>0$ is the conditional moment of the beginning of the inflation phase.
We indicate some of the most obvious properties of the solution (62), (63), (64). The estimates given below are very rough. Let us accept the following estimates for the duration of the inflation time $t_{\mbox{inf}}$, and for the constant $\Lambda_{0}$:
$t_{\mbox{inf}}\cong 10^{-37}\mbox{s},\quad H_{0}\cong 10^{39}\mbox{s}^{-1}.$ (65)
In (65) dimension is restored. Then $H_{0}t_{\mbox{inf}}\cong 100$. Let’s take $\varkappa(t_{0})\cong 1/150$. Assume that during the time $t_{\mbox{inf}}$ the function $\varkappa$ changes insignificantly. Then for $t_{0}<t<t_{0}+t_{\mbox{inf}}$ the solutions (62), (63), (64) take the form
$H(t)\cong H_{0},\quad\tilde{\varepsilon}(t)\cong-\tilde{p}\cong-\Lambda_{0}+3H_{0}^{2}.$ (66)
Thus, during inflation, the scale factor $a(t)$ increased by $(\mathop{\rm exp}\nolimits H_{0}t_{\mbox{inf}})\approx(\mathop{\rm exp}\nolimits 100)\approx 10^{43}$ times.
Assume that when $t>t_{0}+t_{\mbox{inf}}$, the function $\varkappa(t)$ becomes equal to $\varkappa=4/3$. In this case, the solutions (62), (63), (64) give a power-law expansion:
$H(t)\cong\frac{1}{2t},\quad\tilde{\varepsilon}(t)\cong-\Lambda_{0}+\frac{3}{4t^{2}},\quad\tilde{p}\cong\Lambda_{0}+\frac{1}{4t^{2}}.$ (67)
Solution (67) shows that the scale factor and the density of real matter change according to the well-known law, as well as the correct equation of state in the case of ultrarelativistic matter (in (68) dimension restored):
$a(t)\propto\sqrt{t},\quad\rho_{\mbox{real}}=\frac{3}{32\pi Gt^{2}},\quad p_{\mbox{real}}=\frac{1}{3}\varepsilon_{\mbox{real}}.$ (68)
From the equalities (63) and (64) it is clear that as we approach the point of the Big Bang, the quantities $|\tilde{\varepsilon}|$ and $\tilde{p}$ decrease and become equal to
$\tilde{\varepsilon}=-(\Lambda_{0}-3H_{0}^{2}),\quad\tilde{p}=\Lambda_{0}-3(1-\varkappa(t_{0}))H_{0}^{2}\quad\mbox{for}\quad t=t_{0}.$
If we assume that at $t\longrightarrow 0$ we have $\varkappa\longrightarrow 0$ and $3H_{0}^{2}\longrightarrow\Lambda_{0}$, then the energy-momentum tensor tend to zero at the point of the Big Bang. However, it was shown above that such a continuation of the solution to the Big Bang point is impossible due to the development of quantum fluctuations. Nevertheless, the tendency towards an absolute decrease in the energy-momentum tensor in the classical solution when approaching the Big Bang point shows the consistency of the presented classical and quantum approaches.
IV.2 A brief review of the divergent cosmological constant problem
The problem with the cosmological constant is that the energy density of quantum zero-point fluctuations of the vacuum diverges as the fourth power of the cutoff parameter, and there is currently no generally accepted solution to compensate for this enormous energy density.
It seems to us that in this Section above a possible solution to the problem is presented, which is correct in the case when space-time has the property of granularity (lattice) on an ultra-small scale. Indeed, the introduction of a finite bare cosmological constant leads to a completely reasonable solution to Einstein’s equations: in the initial phase we have an exponential expansion of the Universe (inflation regime), which turns into a known expansion according to a power law in the regime of ultrarelativistic matter.
We consider it appropriate to give here a very brief and incomplete overview of attempts to solve the identified problem within the framework of traditional quantum field theory.
In the fundamental review Weinberg (1989) the following facts were stated regarding the problem of divergent energy (divergent cosmological constant) of the ground state in quantum field theory: (i) In flat Minkowski space-time, these divergences, generally speaking, take place, but in the case of supersymmetric theories, the energies of the ground states strictly vanish. (ii) In curved space-time, even in the case of supergravity, the cosmological constant diverges. (iii) The superstring theory does not save the situation either.
Of the later and specialized works, we note the works Krotov and Polyakov (2011); Polyakov (2012); Akhmedov (2014); Akhmedov et al. (2019); Akhmedov (2021); Kamenshchik et al. (2022). In these papers, efforts were made to solve the problem of the cosmological constant in detail, that is, through microscopic analysis. In particular, the probability of the following process was calculated. Let there be a massive particle in the de Sitter space. This particle gives rise to a pair of the same particles for a sufficiently long period of time. This problem has been studied for both free and interacting fields. A similar statement of the problem for massive charged particles in the case of a flat space-time in the presence of a constant electric field leads to the creation of particle-antiparticle pair that weaken the initial electric field. In the case of de Sitter space, pair production also leads to a decrease in the cosmological constant with time. Unfortunately, in these works there is no study of the reverse influence of quantized material fields on the space-time geometry. It is possible that continued efforts in this direction will lead to a solution to the problem of the cosmological constant.
In the paper Zel’Dovich and Starobinskiǐ (1972) the mean of the energy-momentum tensor of a quantized scalar field is calculated in the case of an anisotropic metric, which is considered to be classical and variable in time. Regularization is carried out in the usual way: the vacuum expectation value of the energy-momentum tensor, calculated in the case of a stationary vacuum, is subtracted from the obtained value.
The authors of the paper Kamenshchik et al. (2018) study such models of field theory which, although not supersymmetric, have the same number of boson and fermion degrees of freedom. In this case, the divergences of the highest, fourth degree are eliminated in the quantum mean of the energy-momentum tensor. It is shown what conditions the already renormalized field masses must satisfy in order to reduce all other divergences. This approach can be called a method of fine-tuning the theory, as a result of which the infinite vacuum energy disappears.
The work Appleby and Linder (2020) seems to us to be interesting and complementary to the present work, since a bare cosmological constant is also introduced in Appleby and Linder (2020), and the reduction of the huge vacuum energy is a dynamic effect, not a fine-tuning effect.
Another interesting approach to solving the problem, using the macroscopic thermodynamic ideology, is presented in Klinkhamer and Volovik (2022) (see there also references for the articles of F.R. Klinkhamer and G.E. Volovik). The main idea of this approach is as follows. If the system comes to a state of thermodynamic equilibrium, then a large thermodynamic potential is of interest. Let $\Omega$ be a large thermodynamic potential for the spatial volume $V$. It is known that
$\Omega(\beta,\mu,V)=-P(T,\mu)V.$ (69)
In the case of thermal equilibrium (if it exists), the pressure on the right side of the equality (69) tends to zero, since there is no external pressure at all. Further, the effective energy-momentum tensor of matter is formed by the potential (69). Therefore, the effective energy density of matter, including the vacuum energy, under the condition of thermal equilibrium is estimated as $\varepsilon\sim\Omega/V\longrightarrow 0$. Thus the problem of the divergent cosmological constant is removed.
In the work Wang et al. (2017) the use of the metric (51) is criticized due to the strong fluctuations of all fields on the Planck scales of frequencies and wavelengths. This conclusion is made on the basis of studying the correlators of the energy-momentum tensor components constructed from quantized matter fields. The calculation shows that even in the case of free quantized fields, the vacuum average values of the energy-momentum tensor components and their fluctuations are of the same order and diverge as the fourth power of the cutoff momentum. The fact that the divergences of these quantities are of the fourth order means that these divergences saturate on the scale of the cutoff momentum. In other words, exclusively high-frequency degrees of freedom contribute to the discussed averages. Therefore, we draw the following conclusion: if we describe the dynamics of any degrees of freedom that have frequencies and wave vectors much less than Planck’s, then the high-frequency fluctuations discussed in Wang et al. (2017) do not play a role in such dynamics. Indeed, the correlators between the low-frequency and high-frequency degrees of freedom are in fact equal to zero. This justifies the use of the metric (51) in the Einstein equations discussed above, as well as the assumption that the energy-momentum tensor is uniform in space. However, these assumptions become incorrect when the scale factor and mean of the energy-momentum tensor components become small and the temperature becomes high enough (see Section III).
V Thermodynamic considerations
The purpose of this consideration is to (at least superficially) explain the state equation (61).
The estimation (65) means that
$H_{0}^{2}\ll\Lambda_{0}.$ (70)
It can be seen from Eq. (63) that the maximum frequencies of the degrees of freedom of matter in the modern era are of the order of
$|\omega_{\mbox{max}}|\sim\sqrt{\Lambda_{0}}.$ (71)
Consider the time interval $\Delta t\lesssim H^{-1}$, for which we have
$\Delta a/a\sim H\Delta t\lesssim 1,\quad\Delta t|\omega_{\mbox{max}}|\gg 1.$ (72)
Taking into account Eq. (72), we can assume that for a time interval $\Delta t$ the thermodynamic equilibrium of the vacuum degrees of freedom is realized. This assumption cannot be extended to those degrees of freedom whose frequencies $\Delta t|\omega|\lesssim 1$. But such degrees of freedom make a small contribution to the total energy-momentum tensor. Thus, the degrees of freedom with frequencies in the interval
$H_{0}\sim\Delta t^{-1}\ll|\omega|<|\omega_{\mbox{max}}|$ (73)
are in a state of thermal equilibrium or close to it.
The conclusion about the thermalization of vacuum degrees of freedom in the indicated frequency range (73) can be supported by the following reasoning. There is no indication that the system under study is integrable. Therefore, the dynamics of all distribution functions of the system must be described by the Boltzmann equation. The Boltzmann equation brings the system to a locally equilibrium state in time intervals commensurate with the inverse values of the characteristic frequencies of the dynamic degrees of freedom. This statement is true under the condition of an insignificant change in external conditions during the process of establishing equilibrium. To fulfill the last condition, it is necessary to satisfy the estimates (72) and (73).
When passing to the Euclidean signature by Wick’s rotation $\Delta t=-i\Delta\tau$ , the parameter
$T_{\mbox{vac}}\equiv\beta^{-1}=\hbar(\Delta\tau)^{-1}\approx\hbar H$ (74)
acquires the meaning of local vacuum temperature. Here the dimension is restored. An exact calculation of the local vacuum temperature in de Sitter spacetime has been done in Volovik (2009, 2024, 2023): $T_{\mbox{vac}}=\hbar H/\pi$. Local temperatures differ from the horizon temperature in de Sitter space-time Gibbons and Hawking (1977), although all the temperatures discussed here are of the same order. As can be seen, simple qualitative reasoning within the framework of quantum field theory leads to a correct estimate of the local vacuum temperature.
Let’s make a remark. The calculation of the local vacuum temperature in the paper Volovik (2009) is tied to the space-time of a particular geometry: the de Sitter space-time. On the contrary, our approach to estimating the local vacuum temperature can be applied to the space-time of any geometry. We need only relatively local homogeneity and isotropy of space. The general conditions for the existence and measurability of the local vacuum temperature are not discussed here.
Here one should cite the classic work of Unruh Unruh (1976), in which the problem of calculating the temperature of, for example, the horizon of a black hole, was connected with the dynamics of quantized fields.
Let us determine the temperature value in Kelvin degrees at the beginning of the inflation process, when, according to some estimates $H_{0}\sim 10^{39}\mbox{s}^{-1}$. Then
$T_{0}\sim\frac{\hbar H_{0}}{k}\sim\left(10^{28}\right)^{\circ}\mbox{K}.$ (75)
Here $k$ is the Boltzmann constant. The temperature estimate (75) is within the known temperature estimates in the initial phase of inflation.
Once again, we note that thermodynamic considerations do not apply to low-frequency degrees of freedom. In particular, ordinary real matter may, generally speaking, not be in a state of thermal equilibrium.
According to Eqs. (62) and (74) we have:
$\hbar\mathop{\rm d}\nolimits\beta\sim\mathop{\rm d}\nolimits(1/H)=3/2\varkappa\mathop{\rm d}\nolimits t.$
But in the inflation phase $a(t)=a_{0}e^{Ht}$, and so $\mathop{\rm d}\nolimits t=H^{-1}\mathop{\rm d}\nolimits a/a$. Thus we have:
$\mathop{\rm d}\nolimits\beta/\beta\sim(3/2)\varkappa\mathop{\rm d}\nolimits a/a.$ (76)
Since the temperature decreases in the inflation phase, it can be seen from (76) that $\varkappa(t_{0})>0$.
It can be seen from the first Eq. (59) that the constant $\Lambda_{0}$ cancels out the huge negative energy of the vacuum, so that in the era of power-law expansion only the relatively extremely small positive energy density of real matter affects the dynamics. From the given solution of Einstein’s equations, it can be seen that the huge value of pressure is also mainly reduced by the constant $\Lambda_{0}$. In the presented solution we have $\tilde{p}\sim-\tilde{\varepsilon}\sim\Lambda_{0}$. Such a ratio of energy and pressure is possible only in the general theory of relativity.
Let us introduce the notation $\tilde{\epsilon}_{\mbox{vac}}\equiv(\tilde{\varepsilon}+\Lambda_{0})$. With the help of Eqs. (58) and (60) we find:
$\tilde{\epsilon}_{\mbox{vac}}\equiv(\tilde{\varepsilon}+\Lambda_{0}),\quad\epsilon_{\mbox{vac}}=\frac{\tilde{\epsilon}_{\mbox{vac}}}{8\pi G}=\frac{3H^{2}}{8\pi G}.$ (77)
The last relation is contained in Volovik (2023). According to (74) the vacuum temperature $T\approx\hbar H$. Therefore, the ratio (77) is rewritten as follows:
$\epsilon_{\mbox{vac}}=\lambda\frac{3T^{2}}{8\pi G\hbar^{2}},\quad\lambda\sim 1.$ (78)
Statement. If we know the energy density as a function of temperature, then the free energy density is
$f_{\mbox{vac}}=-T\int_{0}^{T}\frac{\epsilon_{\mbox{vac}}(T^{\prime})}{(T^{\prime})^{2}}\mathop{\rm d}\nolimits T^{\prime}.$ (79)
Checking the Statement is reduced to checking the fact that the function (79) satisfies the following system of equations:
$s_{\mbox{vac}}=-(\partial/\partial T)f_{\mbox{vac}},\quad f_{\mbox{vac}}=\epsilon_{\mbox{vac}}-Ts_{\mbox{vac}}.$ (80)
Here $s_{\mbox{vac}}$ is the local vacuum entropy density. In our case (78) we obtain:
$f_{\mbox{vac}}=-\epsilon_{\mbox{vac}}=-\lambda\frac{3T^{2}}{8\pi G\hbar^{2}},\quad s_{\mbox{vac}}=\lambda\frac{3T}{4\pi G\hbar^{2}}.$ (81)
A more accurate calculation for the quantities (78) and (81) in de Sitter space-time is given in Volovik (2023), which shows that $\lambda=\pi^{2}$.
The estimate $\tilde{p}\cong\Lambda_{0}$ together with the vacuum energy hypothesis $\tilde{\varepsilon}\cong-\Lambda_{0}$ justifies the equation of state (61). In both parts of equality $(\tilde{\varepsilon}+\tilde{p})=\varkappa(\tilde{\varepsilon}+\Lambda_{0})$, the diverging values of the quantities cancel each other out. This fact is the result of solving dynamic equations.
VI Conclussion
In the works Diakonov (2011); Vladimirov and Diakonov (2012); Vladimirov and Diakonov (2014); Volovik (2021) and Vergeles (2021a) the equality
$\left\langle e_{{{\cal V}}_{1}{{\cal V}}_{2}}^{a}\right\rangle=\varkappa^{(0)}_{({{\cal V}}_{1}{{\cal V}}_{2})}\left\langle\Theta^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\right\rangle$ (82)
was proposed for lattice theory of gravity coupled with Dirac field. The relation (82) is justified by the fact that the values under the average transform identically under all symmetry transformations. The relation (50), obtained by direct calculation in the continuum limit at zero temperature, is another argument in favor of the validity of the relation (82).
Thus, if the hypothesis proposed above is correct, then at the maximum possible temperature, the right side of Eq. (82) vanishes according to (49), whence follows
$\left\langle e_{{{\cal V}}_{1}{{\cal V}}_{2}}^{a}\right\rangle_{\mbox{Gauge Fix}}=0.$ (83)
The last equality means that in the lattice theory of gravity in the high-temperature phase, space collapses into a point.
Obviously, in the long-wavelength limit at zero temperature, the contribution of the Dirac field to the energy-momentum tensor is nonzero (50).
The above means that in the lattice theory of gravity there is a phase transition in temperature (possibly more than one). In the high-temperature phase, space collapses into a point, the mean of the energy-momentum tensor is zero and the PT symmetry is not broken. This statement is proved for 4D lattice model of gravity coupled with Dirac field. In the low-temperature phase, these quantities are nonzero and the PT symmetry is broken. The role of the order parameter is played by the mean $\langle e^{a}_{\mu}\rangle$ which becomes nonzero in the low-temperature phase. The process of exponential expansion of space begins, turning into a power law of expansion. During the phase transition from the high-temperature to the low-temperature phase, domains with opposite mean values $\langle e^{a}_{\mu}\rangle$ and $\langle\Theta^{a}_{\mu}\rangle$ can be formed. The domain wall separating such domains was studied in Vergeles (2021b)
Acknowledgements.
I thank Prof. G.E. Volovik for awakening my interest in the thermodynamic study of the problem. I am grateful to Prof. E.T. Akhmedov for numerous discussions and advice in the course of work. This work was carried out as a part of the State Program 0033-2019-0005.